### v2

# Maximum Likelihood Estimates of Parameters in Generalized Gamma Distribution with SeLF Algorithm

Abstract
This undergraduate thesis focuses on the problem of calculating maximum likelihood estimates (MLEs) of parameters in the generalized Gamma distribution using the SeLF algorithm. As an extension of the Gamma distribution, the generalized Gamma distribution can better fit real data and has been widely applied. The research begins by exploring the definition of the generalized Gamma distribution and its similarities and differences from the traditional Gamma distribution. Then, the SeLF and US algorithms are discussed in detail. The SeLF algorithm is a new algorithm based on the Minorization-Maximization (MM) algorithm, which can obtain the local optimal solution with few iterations, with the advantages of fast computation, high accuracy, and good convergence. The US algorithm is a method for finding the zeros of a function, which stands at a higher level than the SeLF algorithm and can improve the convergence speed and stability. This thesis proposes a method for calculating maximum likelihood estimates of the parameters in the generalized Gamma distribution using the SeLF and US algorithms, and presents the practical implementation of the algorithms, as well as simulations and data analysis to evaluate the performance of the proposed methods. The results demonstrate that the SeLF algorithm can achieve more stable and accurate estimates of the parameters in the generalized Gamma distribution more quickly, compared to traditional Newton’s method, which can be useful in various applications. This thesis provides a comprehensive and in-depth exploration of the generalized Gamma distribution and the SeLF algorithm, and proposes a new method for calculating maximum likelihood estimates of parameters, contributing to the development of statistical methods for parameter estimation in complex models. It also provides a reference for further exploring the theory and application of the generalized Gamma distribution. The proposed method in this thesis has important practical significance and application value for solving practical problems.
[Key words]: generalized Gamma distribution; MLE; Second–derivative Lower–bound Function (SeLF) algorithm; Upper-crossing/Solution (US) algorithm
1 Generalized Gamma Distribution
1.1 Introduction
The Gamma distribution is a commonly used probability distribution that has many applications in various fields, such as engineering, physics, and finance. It is a continuous probability distribution that is widely used to model waiting times, survival analysis, and lifetime data. However, there are situations where the Gamma distribution may not be an appropriate model due to its restrictive shape. In the paper [stacy1962generalization], [stacy1962generalization] has introduced a generalization of the Gamma distribution called the generalized Gamma distribution (GGD), which allows for more flexibility in the shape of the distribution. In this section, we discuss the properties and applications of this distribution and compare it to the (traditional) Gamma distribution.
1.2 Definition of Generalized Gamma Distribution
The generalized Gamma distribution is a three-parameter continuous probability distribution that has a flexible shape and can model a wider range of data than the Gamma distribution.
Definition 1 (Generalized Gamma Distribution).
The generalized Gamma distribution is a distribution with the general formula of the probability density function (PDF) given by:
$f(x;\alpha,\beta,\gamma)=\frac{\beta^{\alpha}\gamma}{\Gamma(\alpha/\gamma)}x^{\alpha-1}e^{-(\beta x)^{\gamma}},\ x>0$
where $\alpha,\beta>0$ are the shape parameters, $\gamma>0$ is the scale parameter and $\Gamma$ is the gamma function which is defined by $\Gamma(x)=\int_{0}^{\infty}t^{x-1}e^{-t}{\rm d}t$ for any $x>0$.
The Gamma distribution is a special case of the generalized Gamma distribution when $\gamma=1$. When $\gamma>1$, the generalized Gamma distribution has a heavier tail than the Gamma distribution, while when $\gamma<1$, it has a lighter tail.
1.3 Properties of Generalized Gamma Distribution
The generalized Gamma distribution has several properties that make it a useful model for a wide range of data. Some of these properties are:
  • Flexibility: The generalized Gamma distribution can model a wide range of data shapes due to its flexibility. The shape of the distribution can be adjusted by varying the shape and scale parameters.
  • Moment Generating Function: The moment generating function (MGF) of the generalized Gamma distribution is given by:
    $M(t)=\sum_{n=0}^{\infty}\frac{(\frac{t}{\beta})^{n}}{n!}\cdot\frac{\Gamma(\frac{\alpha+n}{\gamma})}{\Gamma(\frac{\alpha}{\gamma})}$
    where $\Gamma$ denotes the Gamma function. This can be used to derive the moments of the distribution.
  • Cumulative Distribution Function: The cumulative distribution function (CDF) of the generalized Gamma distribution does not have a closed-form solution. However, it can be approximated using numerical methods.
1.4 Applications of Generalized Gamma Distribution
The generalized Gamma distribution has several applications in various fields. Some of these applications are:
  • Lifetime Data: The generalized Gamma distribution can be used to model lifetime data, such as the time to failure of a component in a machine. For a more detailed discussion, please refer to [shanker2016modeling]’s paper [shanker2016modeling].
  • Wind Speed: The generalized Gamma distribution can be used to model wind speed data, which is important in the design of wind turbines. For a more detailed discussion, please refer to the paper [mert2015statistical] by [mert2015statistical].
  • Finance: The generalized Gamma distribution can be used to model the distribution of stock returns, which can help in the analysis of financial data. For a more detailed discussion, please refer to [gomes2006four]’s paper [gomes2006four].
1.5 Comparison with Gamma Distribution
The gamma distribution is a well-known distribution used in many fields, including statistics, physics, engineering, and finance. The generalized Gamma distribution (GGD) is a generalization of the gamma distribution that offers more flexibility in modeling real-world data. In this subsection of the thesis, we compare and contrast the two distributions and highlight the advantages of the GGD.
Definition 2 (Gamma Distribution).
The gamma distribution is a family of continuous probability distributions with two parameters, including a shape parameter $\alpha$ and a scale parameter $\beta$. Its probability density function (PDF) is given by:
$f(x;\alpha,\beta)=\frac{\beta^{\alpha}}{\Gamma(\alpha)}x^{\alpha-1}e^{-\beta x},\ x>0$
where $\Gamma(\alpha)$ is the gamma function.
The gamma distribution has many applications, including in reliability analysis, queuing theory, and finance. However, it is limited in its ability to model data with different shapes and tail behavior. In particular, the gamma distribution has a monotonically decreasing hazard function, which means that the failure rate decreases over time.
The generalized Gamma distribution is a more flexible distribution than the gamma distribution because it allows for different tail behavior and can model data with more complex shapes. The parameter $\gamma$ controls the shape of the distribution, and when $\gamma=1$, the generalized Gamma distribution reduces to the gamma distribution. The generalized Gamma distribution has been used in many applications, including in modeling rainfall, wind speed, and response times in neuroscience.
The gamma distribution is a special case of the generalized Gamma distribution, and therefore, it is less flexible than the generalized Gamma distribution. The generalized Gamma distribution can model data with more complex shapes and tail behavior, making it more suitable for real-world applications. In addition, the generalized Gamma distribution has a more interpretable shape parameter $\gamma$, which allows for easier interpretation of the distribution’s properties. However, the gamma distribution is simpler to work with and has a closed-form solution for many statistical problems, making it more computationally efficient.
In summary, the generalized Gamma distribution is a more flexible distribution than the gamma distribution and offers more options for modeling real-world data. However, the gamma distribution is simpler to work with and has closed-form solutions for many statistical problems. The choice between the two distributions depends on the specific application and the goals of the analysis.
2 Second–derivative Lower–bound Function Algorithm
2.1 Introduction
The paper [tian2025second] by [tian2025second] has proposed the second-derivative lower-bound function (SeLF) algorithm, which is a novel maximization method based on the framework of minorization-maximization (MM) algorithms. The SeLF algorithm is a useful tool for calculating the maximum likelihood estimate (MLE) $\hat{\theta}$ of some parameter in a one-dimensional log-likelihood function $l(\theta)$ (or maximize a more generalized target function). In order to introduce the SeLF algorithm more clearly, we have to spend some space introducing MM algorithms first, because the SeLF algorithm is based on the fundamental principles of MM algorithms.
2.2 Minorization-Maximization Algorithms
Minorization-Maximization (MM) algorithms are a family of iterative optimization algorithms that are commonly used in machine learning and statistical inference. These algorithms are particularly useful for solving optimization problems that involve log-concave likelihood functions or other functions that are difficult to optimize directly. In this subsection, we will provide a formal definition of MM algorithms and discuss some of their key properties and applications. [hunter2004tutorial] compiled the general principles of MM algorithms in their paper [hunter2004tutorial].
2.2.1 Definition
The basic idea behind MM algorithms is to construct a sequence of simpler functions that “minorize” the objective function of interest, and then optimize each of these simpler functions in turn. At each iteration, the algorithm chooses a minorizing function that is a lower bound on the objective function, and then optimizes this minorizing function. This optimization step is known as the “maximization” step, since we are maximizing a lower bound on the objective function.
More formally, suppose we wish to maximize a function $l(\theta)$ over some parameter space $\Theta$. We define a surrogate function $Q(\theta|\theta^{(t)})$ as a function that depends on the current estimate $\theta^{(t)}$, and that satisfies the following two conditions:
  1. 1.
    $Q(\theta|\theta^{(t)})\leq l(\theta)$ for all $\theta\in\Theta$
  2. 2.
    $Q(\theta^{(t)}|\theta^{(t)})=l(\theta^{(t)})$
Condition (1) ensures that $Q(\theta|\theta^{(t)})$ is a lower bound on $l(\theta)$, while condition (2) ensures that the bound is tight at the current estimate $\theta^{(t)}$. Note that there may be many functions that satisfy these conditions; the choice of $Q(\theta|\theta^{(t)})$ will depend on the specific problem being solved. The process of finding the $Q$ function is usually referred to as the minorization step.
Given a minorizing (or surrogate) function $Q(\theta|\theta^{(t)})$, we define the next estimate in the sequence as:
$\theta^{(t+1)}=\arg\max_{\theta\in\Theta}Q(\theta|\theta^{(t)}).$
That is, we choose the value of $\theta$ that maximizes the lower bound $Q(\theta|\theta^{(t)})$. This step is known as the maximization step.
The iteration continues until convergence, typically when the change in the estimate between iterations falls below a predefined tolerance level.
2.2.2 Properties
MM algorithms have several useful properties that make them well-suited for a variety of optimization problems.
Convergence
One of the most important properties of an optimization algorithm is convergence. MM algorithms have been shown to converge under certain conditions. Specifically, if $l(\theta)$ is a concave function and the surrogate function $Q(\theta|\theta^{(t)})$ satisfies the following conditions:
  1. 1.
    $Q(\theta^{(t)}|\theta^{(t)})=l(\theta^{(t)})$
  2. 2.
    $Q(\theta|\theta^{(t)})\leq l(\theta)$ for all $\theta\in\Theta$
  3. 3.
    $\theta^{(t+1)}=\arg\max_{\theta}Q(\theta|\theta^{(t)})$
then the MM algorithm is guaranteed to converge to a stationary point of $l(\theta)$.
Monotonicity
Another important property of optimization algorithms is monotonicity, which means that the objective function value improves or stays the same at each iteration. MM algorithms are not necessarily monotonic since there is another type of minorization-maximization algorithm (usually referred to as generalized minorization-maximization algorithm) which only requires the $Q$ function to be smaller than the $l$ function at any point $\theta\in\Theta$ yet without requiring the $Q$ function and the $l$ function to be equal at point $\theta^{(t)}$, but under certain conditions, they can be monotonic.
Specifically, if the surrogate function $Q(\theta|\theta^{(t)})$ satisfies the following condition:
  • $Q(\theta^{(t)}|\theta^{(t)})=l(\theta^{(t)})$
then the MM algorithm is monotonic, meaning that $l(\theta^{(t+1)})\geq l(\theta^{(t)})$ for all $t$.
Stability
Stability is another important property of optimization algorithms. A stable algorithm is one that does not produce wildly varying solutions for small changes in the input or parameters. MM algorithms are generally stable, but they can exhibit instability if the surrogate function is not well-designed.
To ensure stability, the surrogate function $Q(\theta|\theta^{(t)})$ should satisfy the following condition:
  • $|Q(\theta|\theta^{(t)})-Q(\theta|\theta^{(t^{\prime})})|\leq C|\theta^{(t)}-\theta^{(t^{\prime})}|$
where $C$ is a constant that depends on the choice of surrogate function and $\theta^{(t^{\prime})}$ is a nearby point to $\theta^{(t)}$. This condition ensures that small changes in the input or parameters result in small changes in the surrogate function.
In summary, minorization-maximization (MM) algorithms are a class of iterative optimization methods that are useful for maximizing difficult objective functions. They are guaranteed to converge to a stationary point under certain conditions and can be monotonic under additional conditions. Furthermore, they are generally stable, but stability depends on the choice of surrogate function.
2.3 SeLF algorithm
2.3.1 Definition
The paper [tian2025second] proposes a novel minorization-maximization (MM) method known as the second-derivative lower-bound function (SeLF) algorithm, which is a general approach for iteratively computing the maximum likelihood estimate (MLE) $\hat{\theta}$ of the parameter $\theta$ in a one-dimensional target function of interest, typically the marginal likelihood function $l(\theta)$.
The SeLF algorithm consists of two steps per iteration: a second-derivative lower-bound function step (SeLF-step), which is a special type of minorization step, and a maximization step (M-step).
Let $l(\theta)$ be a one-dimensional target function of interest, such as the log-likelihood function of the parameter $\theta$, and assume that the first-order derivative $l^{\prime}(\theta)$ and the second-order derivative $l^{\prime\prime}(\theta)$ exist. We seek to obtain the MLE of $\theta$ such that $\hat{\theta}=\arg\max_{\theta\in\Theta}l(\theta)$, where $\Theta\subseteq\mathbb{R}\triangleq(-\infty,\infty)$ is the parameter space.
Suppose that there exists a function $b(\theta)$ where $\theta\in\Theta$ such that $l^{\prime\prime}(\theta)\geq b(\theta)$ for all $\theta\in\Theta$, where $b(\theta)$ is called the second-derivative lower-bound function (SeLF) for $l(\theta)$. To construct a minorizing function based on the SeLF $b(\theta)$, we first define
$B(\theta|\theta^{(t)})\triangleq\int_{\theta^{(t)}}^{\theta}b(z){\rm d}z$
and then we introduce the following auxiliary function
$h(\theta|\theta^{(t)})\triangleq l(\theta)-\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z.$
As a result,
$h^{\prime}(\theta|\theta^{(t)})=$ $l^{\prime}(\theta)-B(\theta|\theta^{(t)}),$
$h^{\prime\prime}(\theta|\theta^{(t)})=$ $l^{\prime\prime}(\theta)-b(\theta)\geq 0,$
$h(\theta^{(t)}|\theta^{(t)})=$ $l(\theta^{(t)}),$
$h^{\prime}(\theta^{(t)}|\theta^{(t)})=$ $l^{\prime}(\theta^{(t)})-B(\theta^{(t)}|\theta^{(t)})=l^{\prime}(\theta^{(t)}).$
Since $h^{\prime\prime}(\theta|\theta^{(t)})\geq 0$, indicating that $h(\theta)$ is convex, by the Taylor formula from analysis, we can see that for some specific $\theta^{*}$ between $\theta$ and $\theta^{(t)}$
$h(\theta|\theta^{(t)})$ $=h(\theta^{(t)}|\theta^{(t)})+h^{\prime}(\theta^{(t)}|\theta^{(t)})(\theta-\theta^{(t)})+\frac{1}{2}h^{\prime\prime}(\theta^{*})(\theta-\theta^{*})^{2}$
$\geq h(\theta^{(t)}|\theta^{(t)})+h^{\prime}(\theta^{(t)}|\theta^{(t)})(\theta-\theta^{(t)})$
where the equality holds iff $\theta=\theta^{(t)}$. Therefore, we obtain
$l(\theta)$ $=h(\theta|\theta^{(t)})+\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z$
$\geq h(\theta^{(t)}|\theta^{(t)})+h^{\prime}(\theta^{(t)}|\theta^{(t)})(\theta-\theta^{(t)})+\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z$
$=l(\theta^{(t)})+l^{\prime}(\theta^{(t)})(\theta-\theta^{(t)})+\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z$
$=l^{\prime}(\theta^{(t)})\theta+\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z+c^{(t)}$
which minorizes $l(\theta)$ at $\theta=\theta^{(t)}$ where $c^{(t)}=l(\theta^{(t)})-l^{\prime}(\theta^{(t)})\theta^{(t)}$ is a constant free from $\theta$. So we can define $Q(\theta|\theta^{(t)})\triangleq l^{\prime}(\theta^{(t)})\theta+\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z+c^{(t)}$ as the surrogate function of the MM algorithm.
According to the M-step of minorization-maximization algorithm, letting ${\rm d}Q(\theta|\theta^{(t)})/{\rm d}\theta=0$, the maximizer of the $Q(\theta|\theta^{(t)})$ function is
$\theta^{(t+1)}$ $=\arg\max_{\theta\in\Theta}Q(\theta|\theta^{(t)})$
$=\arg\max_{\theta\in\Theta}\left[l^{\prime}(\theta^{(t)})\theta+\int_{\theta^{(t)}}^{\theta}B(z|\theta^{(t)}){\rm d}z\right]$
$={\rm sol}\left\{\theta\in\Theta:l^{\prime}(\theta^{(t)})+B(\theta|\theta^{(t)})=0\right\}$
$={\rm sol}\left\{\theta\in\Theta:l^{\prime}(\theta^{(t)})+\int_{\theta^{(t)}}^{\theta}b(z){\rm d}z=0\right\}$
where $\theta^{(t)}\in\Theta$.
2.3.2 Properties
The SeLF algorithm has several useful properties. It not only inherits all the excellent characteristics of the MM algorithm but also possesses more specific convergence and stability.
Several important properties are described by the paper [tian2025second] as follows.
Convergence
Definition 3 (Weakly Stable Convergence[tian2025second]).
A sequence $\{\theta^{(t)}\}_{t=0}^{\infty}$ is said to weakly stably converge to a fixed point $\theta^{*}$ if the distance between $\theta^{(t)}$ and $\theta^{*}$ decreases at each iteration, i.e., $|\theta^{(t+1)}-\theta^{*}|<|\theta^{(t)}-\theta^{*}|\ \forall\ t\in\mathbb{N}$.
Definition 4 (Strongly Stable Convergence[tian2025second]).
A sequence $\{\theta^{(t)}\}_{t=0}^{\infty}$ is said to strongly stably converge to a fixed point $\theta^{*}$ if it approaches $\theta^{*}$ in a strictly monotonically increasing or decreasing manner, i.e., $\theta^{(0)}<\theta^{(1)}<\cdots<\theta^{(t)}<\cdots\leq\theta^{*}$ or $\theta^{*}\leq\cdots<\theta^{(t)}<\cdots<\theta^{(1)}<\theta^{(0)}$.
An algorithm is said to have weakly (or strongly) stable convergence if the sequence $\{\theta^{(t)}\}_{t=0}^{\infty}$ generated by the algorithm weakly (or strongly) stably converges to a fixed point $\theta^{*}$. It is worth noting that strong stable convergence implies weak stable convergence.
The paper [tian2025second] established the following theorem:
Theorem 1 (Strongly Stable Convergence[tian2025second]).
Let $\hat{\theta}$ denote the MLE of $\theta$, and let $b(\theta)$ be a SeLF for $l(\theta)$ in $\Theta$. Then, for any initial value $\theta^{(0)}\in\Theta$, the sequence $\{\theta^{(t+1)}\}_{t=0}^{\infty}$ generated by the SeLF algorithm strongly stably converges to $\hat{\theta}$.
Stability
The proof of Theorem 1 in the paper [tian2025second] also guarantees the stability property.
In fact, Theorem 1 establishes that the SeLF algorithm possesses two desirable properties: (i) strongly stable convergence to the MLE $\hat{\theta}$; (ii) independence from any initial values in $\Theta$, which sets it apart from the Newton algorithm.
3 Upper-crossing/Solution Algorithm
3.1 Introduction
In [li2022upper]’s recent work [li2022upper], [li2022upper] introduces a new and versatile root-finding method, named the Upper-crossing/Solution (US) algorithm, which belongs to the class of non-bracketing (or open domain) methods. The US algorithm offers a general approach to iteratively determine the unique root $\theta^{*}$ of a non-linear equation $g(\theta)=0$. Each iteration of the US algorithm consists of two steps: an upper-crossing step (U-step) and a solution step (S-step). During the U-step, the algorithm seeks an upper-crossing function or a $U$-function, denoted by $U(\theta|\theta^{(t)})$, which takes a particular form depending on the $t$-th iteration $\theta^{(t)}$ of $\theta^{*}$, and is derived using a novel concept called the changing direction inequality. In the subsequent S-step, the US algorithm solves the simple $U$-equation, $U(\theta|\theta^{(t)})=0$, to obtain the explicit solution $\theta^{(t+1)}$.
3.2 Definition
Definition 5 (Changing Direction (CD) Inequalities[li2022upper]).
For two functions $h_{1}(x)$ and $h_{2}(x)$ with the same domain $\mathbb{X}$,
$h_{1}(x)\overset{{\rm sgn}(a)}{\geq}h_{2}(x)\ \Longleftrightarrow\ \begin{cases}h_{1}(x)\geq h_{2}(x),&{\rm if}\ a>0,\\ h_{1}(x)=h_{2}(x),&{\rm if}\ a=0,\\ h_{1}(x)\leq h_{2}(x),&{\rm if}\ a<0,\end{cases}$
and
$h_{1}(x)\overset{{\rm sgn}(a)}{\leq}h_{2}(x)\ \Longleftrightarrow\ \begin{cases}h_{1}(x)\leq h_{2}(x),&{\rm if}\ a>0,\\ h_{1}(x)=h_{2}(x),&{\rm if}\ a=0,\\ h_{1}(x)\geq h_{2}(x),&{\rm if}\ a<0.\end{cases}$
Suppose that solving directly for the unique root $\theta^{*}$ of the non-linear equation $g(\theta)=0,\ \theta\in\Theta\subset\mathbb{R}$ is extremely challenging. Without loss of generality, we assume that
$g(\theta)>0,\ \forall\ \theta<\theta^{*}\ {\rm and}\ g(\theta)<0,\ \forall\ \theta>\theta^{*}.$
Alternatively, we can simplify the assumption with CD inequalities as $g(\theta)\overset{{\rm sgn}(\theta-\theta^{*})}{\leq}0$ for all $\theta\in\Theta$. If $g(\theta)<0$ when $\theta<\theta^{*}$, then we can multiply both sides of the nonlinear equation $g(\theta)=0,\ \theta\in\Theta\subset\mathbb{R}$ by 1-1 to obtain a new equation $g(\theta)=0$, which satisfies the above assumption.
The idea behind the US algorithm is as follows: Let $\theta^{(t)}$ denote the $t$-th iteration of $\theta^{*}$ and $U(\theta|\theta^{(t)})$ denote a real-valued function of $\theta$ whose form depends on $\theta^{(t)}$. The function $U(\theta|\theta^{(t)})$ is called a $U$-function for $g(\theta)$ at $\theta=\theta^{(t)}$ if
$g(\theta)\overset{{\rm sgn}(\theta-\theta^{(t)})}{\geq}U(\theta|\theta^{(t)}),\ \forall\ \theta\in\Theta$
where $\theta^{(t)}\in\Theta$.
The conditions indicate that the $U(\theta|\theta^{(t)})$ function dominates the objective function $g(\theta)$ at $\theta=\theta^{(t)}$ for all $\theta\leq\theta^{(t)}$ and the $U(\theta|\theta^{(t)})$ function underestimates $g(\theta)$ at $\theta=\theta^{(t)}$ for all $\theta\geq\theta^{(t)}$. Using the definition of the $U$-function, we conclude that $g(\theta)$ is a $U$-function for 0 at $\theta=\theta^{*}$.
Some basic properties of $U$-functions are given by the paper [li2022upper] as follows.
Proposition 1 (Basic Properties of $U$-functions[li2022upper]).
  1. (a)
    Let $g(\theta)$ be defined in $\Theta$, then $g(\theta)$ is a $U$-function for itself at any interior of $\Theta$;
  2. (b)
    If $U_{1}(\theta|\theta^{(t)})$ and $U_{2}(\theta|\theta^{(t)})$ are two $U$-functions for $g(\theta)$ at $\theta=\theta^{(t)}$, then $U_{1}(\theta|\theta^{(t)})+U_{2}(\theta|\theta^{(t)})]/2$ is also a $U$-function for $g(\theta)$ at $\theta=\theta^{(t)}$;
  3. (c)
    If $g_{1}(\theta)$ and $g_{2}(\theta)$ are two functions defined on $\Theta$ and $U_{1}(\theta|\theta^{(t)})$ is a $U$-function for $g_{1}(\theta)$ at $\theta=\theta^{(t)}$, then $U_{1}(\theta|\theta^{(t)})+g_{2}(\theta)$ is a $U$-function for $g_{1}(\theta)+g_{2}(\theta)$ at $\theta=\theta^{(t)}$.
If a $U$-function can be found, one needs only solve the simple $U$-equation: $U(\theta|\theta^{(t)})=0$ to obtain its solution $\theta^{(t+1)}$. This is denoted as
$\theta^{(t+1)}={\rm sol}\{\theta\in\Theta:U(\theta|\theta^{(t)})=0\}$
where $\theta^{(t)}\in\Theta$.
Therefore, the US algorithm is a general principle for iteratively seeking the root $\theta^{*}$ with strongly stable convergence. Each iteration consists of an upper-crossing step (U-step) and a solution step (S-step), where
U-step: